Fermi National Laboratory
 

Accelerator Update for May 31, 2006 - June 1, 2006

 

Wednesday May 31

The day shift began with Booster, the Antiproton Source (Pbar), The Main Injector (MI), Recycler, and with the Tevatron (TeV) all prepared for accesses to conduct various work. TeV technicians began replacing the P3 magnet.

Cryo system technicians reported at 8:39 AM that TeV sector F4 was free of liquid helium and ready for the magnet change.

Booster came out of access. Operations established beam in the Booster at 1:56 PM.

MI came out of access at 2:52 PM.

NuMI began power supply testing at 3:49 PM.

The TeV received permission to run beam at 5:37 PM.

TeV was out of access by 7:57 PM.

Operations put the TeV separators in conditioning mode at 8:20 PM.

MI experts established beam to the MI dump at 8:36 PM.

TeV turn on began around 8:48 PM.

MI experts established a 150 GeV ramp and then went into a study period at 9:25 PM.

The TeV went into a dry squeeze at 11:47 PM.

Operations reported at 11:57 PM that an Accumulator kicker (A:IKIK) stepper motor controller had failed. Operators looked for a spare card.

Thursday June 1

The midnight shift began with Pbar studies, with the NuMI turn on progressing, and with TeV at 980 GeV.

Operations reversed the polarities for the TeV separators at 2:15 AM.

A Recycler expert began a study at 3:18 AM.

The TeV came out of its dry squeeze at 5:44 AM and decelerated to 150 GeV.

Operations began preparing the TeV for an access at 8 AM to remove a vacuum pump at sector B1.

The Plans for Today

The plan is to get out of the TeV and establish beam. We also want to get beam to NuMI. Accesses to the TeV will now become very limited.

Machine Reports

Linac

Okay. The Neutron Therapy Facility will treat patients today.

Booster

Techs repaired the vertical IPM and a LCW leak on the Q800 magnet.

MI

Experts are ready to send 150 GeV beam to the TeV and 120 GeV beam to NuMI. They are ready to support all other machines.

Antiproton Source

Techs repaired the Debuncher Bend buss power supply. Experts conducted Accumulator aperture scans. Experts want to go to the 120 GeV mode this afternoon.

TeV

All the power supplies have been on and separator conditioning continues. The B11 separator has been vacuum leak checked and tested out okay. The E11 vertical flying wire was fixed. They’ve installed the sector F43 dipole and also repaired a F43 LCW leak.

Recycler

Experts continue to conduct proton studies. They also tested their BPMs and multiwires. The SF6 water skid has been repaired and they hope to have an electron beam by this evening. They are ready for antiprotons.

Switchyard 120

The TeV sector A0 quadrupole magnet had been staged for installation. It will take about 6 hours to install when an access can be arranged.

MiniBooNE

Experts are ready for beam. They have checked their DAQ. (Booster expects to be able to deliver the intensity MiniBooNE needs by next week.)

NuMI

The dehumidifier units, chillers, and special LCW systems work have been completed. Techs are investigating a gas flow problem. Experts want to conduct target scans later today.

Collider Detector at Fermilab

CDF experts closed up their detector. They took cosmic rays and found a system off. They need an access to investigate this further.

D-Zero

The survey crew found that the detector was not completely closed; techs completed the closure. Experts found that some noise had returned; they continued to investigate. Experts want to commission their silicon detectors with no beam.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/28/2007    email Fermilab